mistyrose:

Can someone help me with some questions in graphing?? Please I will post Questions in comments

mistyrose:

Graph f(x)=2x−1 and g(x)=−x+5 on the same coordinate plane. What is the solution to the equation f(x)=g(x) ? Enter your answer in the box. This is one problem

mistyrose:

1 attachment
mistyrose:

Let ​ f(x)=x+2 ​ and ​ g(x)=2x+1 ​ . Graph the functions on the same coordinate plane. What are the solutions to the equation f(x)=g(x) ? Enter your answers in the boxes.

mistyrose:

1 attachment
mistyrose:

x=−1 x = 0 x = 1 x = 2 x = 3 x = 4 x = 5

mistyrose:

1 attachment
Vocaloid:

Graph f(x)=2x−1 and g(x)=−x+5 on the same coordinate plane. What is the solution to the equation f(x)=g(x) ? ^ you just need to graph f(x) and g(x) on the same graph, you can either do this by hand or use software (wolframalpha.com or desmos should work), then find the point where they intersect lmk if you're having trouble or want to check your answer

mistyrose:

Thanks